Circular No. 9036

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 G1 (STEREO) Further to IAUC 9032, K. Battams reports that a slightly diffuse comet

of mag ∼ 10–11 with no apparent tail (discovery observation tabulated below) has been found on SECCHI HI-1B images (cf. IAUC 8930, 8955).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 G1 Apr. 3.572 21^{h}03.^{m}2 − 6^{o}25′ HI* JR 2009-G30

The reduced astrometry by B. G. Marsden, together with the following preliminary parabolic orbital elements (from 36 observations, Apr. 3–6) and an ephemeris, appear on MPEC 2009-G30.

T = 2009 Apr. 12.635 TT  ω = 170.^{o}978
Ω = 120.785 }2000.0
q = 1.13701 AU  i = 107.788

The following geocentric ephemeris assumes H_{10} = 9.0.

2009TT α_{2000} δ_{2000} Δ r ϵ β Mag.
Apr. 9 22^{h}32.^{m}70 − 0^{o}09.^{′}4 1.704 1.139 40.^{o}1 34.^{o}5 10.7
11 22 36.03 − 1 15.1 1.674 1.137 41.5 35.7 10.7
13 22 39.45 − 2 23.3 1.644 1.137 42.9 36.9 10.6
15 22 42.97 − 3 34.1 1.614 1.138 44.3 38.0 10.6
17 22 46.61 − 4 47.9 1.583 1.139 45.8 39.2 10.6
19 22 50.36 − 6 04.7 1.551 1.142 47.3 40.3 10.5

COMETS C/2008 Y6–Y9 (SOHO) Additional Kreutz sungrazers were found on SOHO website images

(cf. IAUC 9032). C/2008 Y6 (mag ∼ 6.5–7), C/2008 Y7 (mag ∼ 6–6.5), and C/2008 Y8 mag ∼ 7–7.5) were small and stellar in appearance in C3 images; C2 images showed C/2008 Y6 and C/2008 Y8 as slightly diffuse, and C/2008 Y7 as teardrop-shaped, with C/2008 Y6 and C/2008 Y7 each showing a short, faint tail. C/2008 Y9 was faint (mag ∼ 8) and diffuse.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 Y6 Dec. 21.138 18^{h}04.^{m}5 −27^{o}12′ C3/2 MU 2009-E63
C/2008 Y7  21.154 18 04.7 −27 19 C3/2 MU 2009-E63
C/2008 Y8  21.571 18 05.8 −25 52 C3/2 MK 2009-E63
C/2008 Y9  21.588 18 04.6 −25 12 C2  MK 2009-E63
2009 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green